# Tapping into the Black Box: Uncovering Aligned Representations in Pretrained Neural Networks

### v1

@@
Abstract
In this paper we argue that ReLU networks learn an implicit linear model we can actually tap into. We describe that alleged model formally and show that we can approximately pull its decision boundary back to the input space with certain simple modification to the backward pass. The resulting gradients (called excitation pullbacks) reveal high-resolution input- and target-specific features of remarkable perceptual alignment on a number of popular ImageNet-pretrained deep architectures. This strongly suggests that neural networks do, in fact, rely on learned interpretable patterns that can be recovered after training. Thus, our findings may have profound implications for knowledge discovery and the development of dependable artificial systems.
Figure 1: A rudimentary 5-step pixel-space gradient ascent guided by excitation pullbacks for pretrained ResNet50. Each cell shows the difference between the perturbed and clean image, targeting the class in the column. Diagonal: original class; off-diagonal: counterfactuals. Last column: randomly selected extra label. See Section 5 for details.
@@ 1 Introduction
This work investigates gradient-based explanations in ReLU networks, with a focus on convolutional architectures commonly used in computer vision. While our empirical results and theoretical developments concentrate on standard ReLU networks trained on image classification tasks, the core arguments are sufficiently general to apply - with the appropriate modifications - to other activation functions and modalities.
We introduce a simple but effective method for interpreting neural predictions via excitation pullbacks, which isolate the robust informative signal in neural activations from the noise introduced by local data peculiarities. We motivate this idea extensively and demonstrate its empirical effectiveness on the following ImageNet-pretrained ReLU models: ResNet50 (He et al., 2015), VGG11_BN (Simonyan and Zisserman, 2015), and DenseNet121 (Huang et al., 2018). These specific architectures were chosen for reasons detailed in Section 5. We also give compelling theoretical arguments backing the claim that excitation pullbacks are indeed faithful local approximations of the network’s decision boundary. In particular, we hypothesize that they reveal a simpler linear model approximated by the trained network, encoded in the pre-activation patterns.
The paper outline is as follows:
  • Section 2 introduces necessary notions and notational conventions; in particular, the gating function and gating-induced pullbacks are described;
  • Section 3 defines tensor field pullbacks and shows that neural networks are linear in their path space;
  • Section 4 defines excitation pullbacks and postulates that they approximate the decision boundary of an implicit kernel machine that is learned in the network’s pre-activation patterns;
  • Section 5 shows that excitation pullbacks of selected architectures pretrained on the ImageNet are perceptually aligned and can be used to produce faithful and meaningful local feature accentuations by just a few steps of gradient ascent along their directions;
  • Section 6 describes the related works, indicates a possible biological feasibility of presented methods and discusses their limitations.
This paper thus bridges practical gradient visualization with theoretical insights into neural function, suggesting a path toward more transparent and robust deep models.
@@ 2.1 ReLU networks
Let $M$ be a feedforward ReLU network with parameters $\theta\in\mathbb{R}^{d_{net}}$, expressible as a composition of affine transformations and pointwise ReLU nonlinearities.111we omit max-pooling layers from the analysis to improve clarity, while noting that similar arguments are likely to apply due to the structural similarity between max-pooling and ReLU; see Section 5.3. Let $L$ denote the number of layers, and let $d_{\ell}$ be the width of the $\ell$-th layer, with $d_{0}$ denoting the input dimension. Let $x_{0}$ denote the identity function on $\mathbb{R}^{d_{0}}$, i.e. $x_{0}(x):=x$. The network applies the following operations:
$z_{\ell}=W_{\ell}x_{\ell-1}+b_{\ell},\quad x_{\ell}=\text{ReLU}(z_{\ell}),\qquad\text{for }\ell=1,\dotsc,L$
where $W_{\ell}\in\mathbb{R}^{d_{\ell}\times d_{\ell-1}}$, $b_{\ell}\in\mathbb{R}^{d_{\ell}}$, and $\text{ReLU}(z)=\max(0,z)$ is applied element-wise. We define the overall network function $M:\mathbb{R}^{d_{0}}\to\mathbb{R}^{d_{L}}$ by the recursive application of the layers, i.e.,
$M(x_{0}):=x_{L}$
@@
where $y_{c}$ is a trainable set of weights $y_{c}\in\mathbb{R}^{d_{L}}$ (a neuron in the network’s head) and $y_{c}^{\top}$ is it’s corresponding covector, i.e. the linear action of $y_{c}$ on $\mathbb{R}^{d_{L}}$ by the dot product: $y_{c}^{\top}x=\langle y_{c},x\rangle$. Standard classification network with $C$ classes has $C$ corresponding vectors $y_{c}$. We will skip the lower index $c$ assuming that it’s fixed.
@@
$\tilde{x}_{\ell-1}=\begin{bmatrix}x_{\ell-1}\\ 1\end{bmatrix}\in\mathbb{R}^{d_{\ell-1}+1},\quad\tilde{W}_{\ell}=\begin{bmatrix}W_{\ell}&b_{\ell}\\ 0\quad\cdots\quad 0&1\end{bmatrix}\in\mathbb{R}^{(d_{\ell}+1)\times(d_{\ell-1}+1)}$
@@
$\tilde{z}_{\ell}=\tilde{W}_{\ell}\tilde{x}_{\ell-1}$
@@
with the understanding that the bias is included in the matrix $W_{\ell}$ and the input vector $x_{\ell-1}$. We will occasionally write $\mathbb{R}^{d^{*}_{0}}\subset\mathbb{R}^{d_{0}}$ to denote the true input space (without the implicitly added coordinate). Notice that, by construction, the $\mathbb{R}^{d^{*}_{0}}$ remains the actual domain of $f$.
@@ 2.2 Gating representation of ReLU networks
where $g_{\ell}\in\{0,1\}^{d_{\ell}}$ is a binary gating vector defined by
@@
Since $z_{\ell}$ is itself a function of the input $x_{0}$, we may define a gating function
$G:\mathbb{R}^{d_{0}}\to\prod_{\ell=1}^{L}\{0,1\}^{d_{\ell}},\quad G(x):=(g_{1},\dots,g_{L})$ (1)
Let $G_{\ell}(x)=\mathrm{diag}(g_{\ell})\in\mathbb{R}^{d_{\ell}\times d_{\ell}}$ denote the diagonal matrix whose diagonal entries are given by $g_{\ell}$. Then, the network output $M(x)=x_{L}$ can be written as a product of (augmented) weight matrices interleaved with input-dependent diagonal matrices:
@@
This representation makes the piecewise-linear structure of ReLU networks explicit: for a fixed input $x$, the gating pattern $(g_{1},\dots,g_{L})\in\prod_{\ell=1}^{L}\mathbb{R}^{d_{\ell}}$ determines a purely linear computation.
@@
$Z:\mathbb{R}^{d_{0}}\to\prod_{\ell=1}^{L}\mathbb{R}^{d_{\ell}},\quad Z(x):=(z_{1},\dots,z_{L})$ (2)
@@ 2.3 Gating-induced networks
where $\Lambda_{\ell}(x)=\mathrm{diag}(\lambda_{\ell}(x))\in\mathbb{R}^{d_{\ell}\times d_{\ell}}$ is the diagonal matrix formed from the $\ell$-th component $\lambda_{\ell}$ of $\Lambda(x)$. In particular, $M$ is $G$-induced, i.e. $M=M_{G}$. Note that $M_{\Lambda}$ implicitly depends on the network parameters $\theta$.
@@ 2.5 Gating-induced pullbacks
$v^{\top}_{\Lambda}:=y^{\top}\vec{M}_{\Lambda}$
We will call both $v^{\top}_{\Lambda}$ and the corresponding vector field $v_{\Lambda}$ a $\Lambda$-pullback. In particular, we have:
$v_{G}^{\top}=y^{\top}\vec{M}_{G}$ (6)
Now, the following sequence of equations can be easily verified:
@@ 2.6 Training dynamics
[…]
3 Neural networks are linear in their path space
@@ 3.1 Network action as atom filtering
Let’s dive deeper into the structure of $M$. Fix a hidden neuron $u_{\ell}[i]$, i.e. the one corresponding to the $i$-th output in the $\ell$-th layer of $M$. Due to the compositional nature of the network, we can define the pullback $v_{\Lambda}|_{(\ell,i)}$ of $u_{\ell}[i]$ in the same way we did for $y$ just consider the appropriate subnetwork below $u_{\ell}[i]$ as the backbone. Moreover, the full-network pullback $v_{\Lambda}$ can be expressed as a weighted sum of the pullbacks of neurons from previous layers, for example:
@@
[…]
@@
Where $(i_{1},\ldots,i_{L})$ is the path through layers $1,\ldots,L$, i.e. the selection of indices $i_{\ell}\in\{1,\ldots,d_{\ell}\}$. Now rearrange the factors:
@@
Let $\mathcal{P}_{k}=\{(p_{k},\ldots,p_{L}):p_{\ell}\in\{k,\ldots,d_{\ell}\}\}$ denote the set of all paths through layers $k,\ldots,L$, fix $p=(p_{1},\ldots,p_{L})$, i.e. $p\in\mathcal{P}_{1}$ and define the path activity function $\Lambda^{p}:\mathbb{R}^{d_{0}}\to[0,1]$ as:
$\Lambda^{p}$ $:=\prod_{\ell=1}^{L}\lambda_{\ell}[p_{\ell}]$ (10)
@@
Note 3.1 Equation 12 expresses $v_{\Lambda}$ as a weighted sum of constant vector fields (we may call them atoms) indexed by $\mathcal{P}_{1}$ with the only dependence on $x$ being via $\Lambda^{p}$. In case of $\Lambda=G$ this amounts to filtering the atoms by a input-specific binary mask.
@@ 3.2 Tensor field pullbacks
Setting $P_{1}=|\mathcal{P}_{1}|=\prod_{\ell=1}^{L}d_{\ell}$ and some enumeration of $\mathcal{P}_{1}$ we may lift $\Lambda^{p}$ to the function $\tilde{\Lambda}:\mathbb{R}^{d_{0}}\to\mathbb{R}^{P_{1}}$ by setting:
@@
hence we will call $\mathbb{R}^{P_{1}}$ a tensor product and refer to it’s elements as tensors. Similarly $\tilde{\Lambda}$ will be called a tensor field over $\mathbb{R}^{d_{0}}$.
@@
It’s easy to see that $v_{\Lambda}=v_{\tilde{\Lambda}}$ and $f_{\Lambda}=f_{\tilde{\Lambda}}$ (by Equations 712 and 13), so we may safely overload the notation.
Note 3.2 $\tilde{\Lambda}(x)\in\bigotimes_{\ell=1}^{L}\mathbb{R}^{d_{\ell}}$ is a rank-1 (i.e. pure) tensor, while $\tau(x)$ is allowed to be any tensor.

### v3

# Pulling Back the Curtain on ReLU Networks

@@
Pulling Back the Curtain on ReLU Networks
Maciej Satkiewicz
314 Foundation, Kraków
maciej.satkiewicz@314.foundation
Since any ReLU network is piecewise affine, its hidden units can be characterized by their pullbacks through the active subnetwork, i.e., by their gradients (up to bias terms). However, gradients of deeper neurons are notoriously misaligned, which obscures the network’s internal representations. We posit that models do align gradients with data, yet this is concealed by the intrinsic noise of the ReLU hard gating. We validate this intuition by applying soft gating in the backward pass only, reducing the local impact of weakly excited neurons. The resulting modified gradients, which we call excitation pullbacks, exhibit striking perceptual alignment on a number of ImageNet-pretrained architectures, while the rudimentary pixel-space gradient ascent quickly produces easily interpretable input- and target-specific features. Inspired by these findings, we formulate the path stability hypothesis, claiming that the binary activation patterns largely stabilize during training and get encoded in the pre-activation distribution of the final model. When true, excitation pullbacks become aligned with the gradients of a kernel machine that mainly determines the network’s decision. This provides a theoretical justification for the apparent faithfulness of the feature attributions based on these pullbacks, potentially even leading to mechanistic interpretability of deeper models. Incidentally, we give a possible explanation for the effectiveness of Batch Normalization and Deep Features, together with a novel perspective on the network’s internal memory and generalization properties. We release the code and an interactive app for easier exploration of the excitation pullbacks.
Figure 1: A rudimentary 5-step pixel-space gradient ascent guided by excitation pullbacks for pretrained ResNet50. Each cell shows the difference between the perturbed and clean image, targeting the class in the column. Diagonal: original class; off-diagonal: counterfactuals. Last column: randomly selected extra label. See Section 6 for details. Best viewed digitally.
@@ 1 Introduction
Understanding how neural networks organize their internal representations remains a longstanding and fundamental challenge in deep learning. In this paper we report the results of a less common line of research, where our primary goal was not incremental leaderboard improvement, but rather breaking through conceptual barriers. We’ve published111github.com/314-Foundation/ResearchLog the weekly research log to facilitate the discussion on this style of scientific inquiry. While we reduce the problem space to ReLU convolutional networks for vision tasks, our methods likely generalize to other activations with similar gating behavior, and potentially offer insights into a broader range of architectures and modalities.
We originally set out by experimenting with prototypical-part based models such as ProtoPNet (Chen et al., 2019). However, this method of constructing self-explainable networks seems to have an unredeemable limitation in that it hinges on the black-box encoder network. To address this, we shifted our focus toward simpler, mechanistically interpretable architectures on CIFAR-10, with the goal of achieving gradient-data alignment. These experiments revealed an intriguing phenomenon: avoiding spatial gradient interference leads to substantially more interpretable gradients. Through a series of ablation studies, we then discovered that it is in fact possible to recover interpretable gradients in standard convolutional architectures by simply modifying the backward pass, leaving the forward computation intact. This led us to introduce what we call excitation pullbacks. The method turned out to scale to widely used ImageNet-pretrained architectures - including non-ReLU ones. However, we intentionally restrict the scope of this work to the following architectures: ResNet50 (He et al., 2015), VGG11_BN (Simonyan and Zisserman, 2015), and DenseNet121 (Huang et al., 2018); for reasons detailed in Section 6.
The striking perceptual alignment of excitation pullbacks strongly suggests a deeper cause. Our primary concern was to validate theoretically that the produced explanations are reasonably faithful to the network’s decision process. Building on our previous investigations we developed a formal framework that allows us to analyze the pullbacks from the path perspective, where path is any sequence of neurons selected from the consecutive layers. In particular, we observe that ReLU networks correspond to linear models in the path space (the Hilbert space of formal linear combinations of all paths), under a feature map given by the tensor product of the binary activation vectors (encoding the on/off states of neurons) of subsequent layers (see Section 3 for details). This made us suspect that these feature maps might largely stabilize during training in the form of a higher-rank tensor field, which we call a path tensor feature (see Hypothesis 1). If that’s indeed the case, then the network behaves like a piecewise linear kernel machine with excitation pullbacks approximating directionally it’s gradient, so that the feature attributions they produce can be considered faithful to the network’s internal circuitry. This would translate to excitation pullbacks constituting a mechanistic interpretation of deep models in a style reminiscent of the “neural circuits” program (Cammarata et al., 2020) - but considerably streamlined by the pullback computation.
Incidentally, the above considerations provide a possible explanation for the widespread effectiveness of Batch Normalization (Ioffe and Szegedy, 2015) in accelerating training, as normalizing neural activations significantly speeds up the stabilization of activation patterns and, consequently, the emergence of the path tensor feature and the global linear structure of the network’s internal representations, see Section 5.3. Furthermore, the success of Deep Features (Zhang et al., 2018) as a perceptual metric could be attributed to its kernel structure and the close relation to the path tensor feature (see Section 5.2). Interestingly, path tensor features allow us to view neural networks as associative memory systems, with excitation pullbacks as their contents (Section 5.4).
Our work demonstrates that standard pretrained networks do in fact align gradients with data, and highlights how one can extract interpretable representations of hidden neurons. Furthermore, the proposed formal framework sheds light on the internal organization of neural networks, and even offers a potential explanation for their generalization capabilities: if the stability condition expressed in Hypothesis 1 is satisfied, then the network’s decisions can be well approximated by a kernel machine whose gradient is feasibly accessible by the excitation pullback. The striking perceptual alignment observed in our experiments strongly suggests that this may indeed be the case, at least to some extent.
The paper is organized as follows:
Section 2 introduces necessary notions and notational conventions; in particular, the gating function and gating-induced pullbacks; Section 3 defines tensor field pullbacks and shows that ReLU networks correspond to linear maps in the path space; Section 4 introduces the excitation pullbacks and discusses the inherent gradient noise; Section 5 discusses the path stability hypothesis and its implications; Section 6 details the pullback visualization method used in the paper; Section 7 describes the related works and discusses the limitations of presented approach.
We release the code222github.com/314-Foundation/ExcitationPullbacks and the interactive demo app333 huggingface.co/spaces/msat/ExcitationPullbacks to make the exploration of excitation pullbacks easily accessible.
@@ 2.1 ReLU networks
Let $M$ be a feedforward ReLU network with parameters $\theta\in\mathbb{R}^{d_{net}}$, expressible as a composition of affine transformations and pointwise ReLU nonlinearities.444we omit max-pooling layers from the analysis to improve clarity, noting that similar arguments are likely to apply due to the structural similarity between max-pooling and ReLU; see Section 6.3. Let $L$ denote the number of layers, and let $d_{\ell}$ be the width of the $\ell$-th layer, with $d_{0}$ denoting the input dimension. Let $x_{0}$ denote the identity function on $\mathbb{R}^{d_{0}}$, i.e. $x_{0}(x):=x$. The network applies the following operations:
$z_{\ell}:=W_{\ell}x_{\ell-1}+b_{\ell},\quad x_{\ell}:=\text{ReLU}(z_{\ell}),\qquad\text{for }\ell=1,\dotsc,L$
where $W_{\ell}\in\mathbb{R}^{d_{\ell}\times d_{\ell-1}}$, $b_{\ell}\in\mathbb{R}^{d_{\ell}}$, and $\text{ReLU}(z)=\max(0,z)$ is applied element-wise. We identify $M$ with the overall network function $x_{L}:\mathbb{R}^{d_{0}}\to\mathbb{R}^{d_{L}}$, i.e., $M(x):=x_{L}(x)$. We’ll often abuse the notation and write $x_{\ell}\in\mathbb{R}^{d_{\ell}}$, meaning the value of $x_{\ell}(x)$ for some fixed $x\in\mathbb{R}^{d_{0}}$; same for $z_{\ell}$ and other similar functions.
@@
where $y_{c}$ is a trainable set of weights $y_{c}\in\mathbb{R}^{d_{L}}$ (a neuron in the network’s head) and $y_{c}^{\top}$ is it’s corresponding covector, i.e. the linear action of $y_{c}$ on $\mathbb{R}^{d_{L}}$ by the dot product: $y_{c}^{\top}x=\langle y_{c},x\rangle$. Standard classification network with $C$ classes has $C$ corresponding vectors $y_{c}$. We will usually skip the lower index $c$ assuming that it’s fixed.
@@
$\tilde{x}_{\ell-1}:=\begin{bmatrix}x_{\ell-1}\\ 1\end{bmatrix}\in\mathbb{R}^{d_{\ell-1}+1},\quad\tilde{W}_{\ell}:=\begin{bmatrix}W_{\ell}&b_{\ell}\\ 0\quad\cdots\quad 0&1\end{bmatrix}\in\mathbb{R}^{(d_{\ell}+1)\times(d_{\ell-1}+1)}$
@@
$\tilde{z}_{\ell}:=\tilde{W}_{\ell}\tilde{x}_{\ell-1}$
@@
with the understanding that the bias is included in the matrix $W_{\ell}$ and the input vector $x_{\ell-1}$. We will occasionally write $\mathbb{R}^{d^{*}_{0}}\subset\mathbb{R}^{d_{0}}$ to denote the true input space (without the implicitly added coordinate). Notice that $\mathbb{R}^{d^{*}_{0}}$ remains the actual domain of $f$, as the extra coordinate is fixed.
@@ 2.2 Gating representation of ReLU networks
where $\odot$ is the Hadamard product and $g_{\ell}\in\{0,1\}^{d_{\ell}}$ is a binary gating vector defined by
@@
Since $g_{\ell}$ is itself a function of the input $x$, we may define the gating function:
$G:\mathbb{R}^{d_{0}}\to\prod_{\ell=1}^{L}\{0,1\}^{d_{\ell}},\quad G(x):=(g_{1}(x),\dots,g_{L}(x))$ (1)
Let $G_{\ell}(x)=\mathrm{diag}(g_{\ell}(x))\in\{0,1\}^{d_{\ell}\times d_{\ell}}$ denote the diagonal matrix whose diagonal entries are given by $g_{\ell}(x)$. Then, the network output can be written as a product of (augmented) weight matrices interleaved with input-dependent diagonal matrices:
@@
This representation makes the piecewise-linear structure of ReLU networks explicit: for a fixed input $x$, the gating pattern $(g_{1},\dots,g_{L})\in\prod_{\ell=1}^{L}\{0,1\}^{d_{\ell}}$ determines a purely linear computation.
@@
$Z:\mathbb{R}^{d_{0}}\to\prod_{\ell=1}^{L}\mathbb{R}^{d_{\ell}},\quad Z(x):=(z_{1}(x),\dots,z_{L}(x))$ (2)
@@ 2.3 Gating-induced networks
where $\Lambda_{\ell}(x)=\mathrm{diag}(\lambda_{\ell}(x))\in[0,1]^{d_{\ell}\times d_{\ell}}$ is the diagonal matrix formed from $\lambda_{\ell}(x)$. In particular, $M$ is $G$-induced, i.e. $M=M_{G}$. Note that $M_{\Lambda}$ implicitly depends on the network parameters $\theta$.
@@ 2.5 Gating-induced pullbacks
$v^{\top}_{\Lambda}:=y^{\top}\vec{M}_{\Lambda}$ (6)
We will call both $v^{\top}_{\Lambda}$ and the corresponding vector field $v_{\Lambda}$ a $\Lambda$-pullback. In particular, we have $v_{G}^{\top}=y^{\top}\vec{M}_{G}$. The following sequence of equations can be easily verified:
@@ 2.6 Training dynamics
Network parameters $\theta$ and $u$ change during training. We index time steps by $t\in\{0,\ldots,T\}$, where $t=0$ denotes the initialization and $t=T$ corresponds to the final state of training. Whenever an upper index $(t)$ appears, it indicates the value of the corresponding object at training step $t$, for example $(y^{(t)},\theta^{(t)})$ denotes the network parameters and $G^{(t)}$ denotes the gating function induced by $M^{(t)}$. For notational brevity, we omit the index $t=T$ when referring to the final trained model.
3 The path space and tensor field pullbacks
@@ 3.1 Network action as atom filtering
Let’s dive deeper into the structure of $M$. Fix a hidden neuron $u_{\ell}[i]$, i.e. the one corresponding to the $i$-th output in the $\ell$-th layer of $M$. Due to the compositional nature of the network, we can define the pullback $v_{\Lambda}|_{(\ell,i)}$ of $u_{\ell}[i]$ in the same way we did for $y$ - just consider the appropriate subnetwork below $u_{\ell}[i]$ as the backbone. Moreover, the full-network pullback $v_{\Lambda}$ can be expressed as a weighted sum of the pullbacks of neurons from previous layers, for example:
@@
i.e. $v_{\Lambda}$ is a weighted sum of layer-$L$ pullbacks. This means that neurons in layer $L$ contribute particular directions $v_{\Lambda}|_{(L,i)}$ to the full pullback $v_{\Lambda}$, weighted by $\lambda_{L}[i_{L}]$, giving an explicit sense to the intuition that neurons form a compositional hierarchy of feature detectors.
@@
Where $(i_{1},\ldots,i_{L})$ is the path through layers $1,\ldots,L$, i.e. the selection of indices $i_{\ell}\in\{1,\ldots,d_{\ell}\}$. Now let’s rearrange the factors:
@@
Let $\mathcal{P}_{k}=\{(p_{k},\ldots,p_{L}):p_{\ell}\in\{k,\ldots,d_{\ell}\}\}$ denote the set of all paths through layers $k,\ldots,L$, fix $p=(p_{1},\ldots,p_{L})\in\mathcal{P}_{1}$ and define the path activity function $\Lambda^{p}:\mathbb{R}^{d_{0}}\to[0,1]$ as:
$\Lambda^{p}(x):=\prod_{\ell=1}^{L}\lambda_{\ell}(x)[p_{\ell}]$ (10)
@@
Note 3.1 Equation 11 expresses $v_{\Lambda}$ as a weighted sum of constant vector fields (atoms) indexed by $\mathcal{P}_{1}$, with the only dependence on $x$ being via $\Lambda^{p}$. In case of $\Lambda=G$ this amounts to filtering the atoms by an input-specific binary mask.
@@ 3.2 Tensor field pullbacks
Setting $P_{1}=|\mathcal{P}_{1}|=\prod_{\ell=1}^{L}d_{\ell}$ and choosing some enumeration of $\mathcal{P}_{1}$ we may lift $\Lambda^{p}$ to the function $\tilde{\Lambda}:\mathbb{R}^{d_{0}}\to\mathbb{R}^{P_{1}}$ as:
@@
We will call $\mathbb{R}^{P_{1}}$ a tensor product and refer to it’s elements as tensors. Similarly, $\tilde{\Lambda}$ will be called a tensor field over $\mathbb{R}^{d_{0}}$.
@@
It’s easy to see that $v_{\Lambda}=v_{\tilde{\Lambda}}$ and $f_{\Lambda}=f_{\tilde{\Lambda}}$ (Equations 711 and 12), so we may safely overload the notation.
Note 3.2 $\tilde{\Lambda}(x)\in\bigotimes_{\ell=1}^{L}\mathbb{R}^{d_{\ell}}$ is a rank-1 tensor, while $\tau(x)$ can have arbitrary rank.